# Diffusion Models in 3D Vision: A Survey

### v1

@@ Abstract
In recent years, 3D vision has become a crucial field within computer vision, powering a wide range of applications such as autonomous driving, robotics, augmented reality (AR), and medical imaging. This field relies on the accurate perception, understanding, and reconstruction of 3D scenes from 2D data sources like images and videos. Diffusion models, originally designed for 2D generative tasks, offer the potential for more flexible, probabilistic approaches that can better capture the variability and uncertainty present in real-world 3D data. However, traditional methods often struggle with efficiency and scalability. In this paper, we review the state-of-the-art approaches that leverage diffusion models for 3D visual tasks, including but not limited to 3D object generation, shape completion, point cloud reconstruction, and scene understanding. We provide an in-depth discussion of the underlying mathematical principles of diffusion models, outlining their forward and reverse processes, as well as the various architectural advancements that enable these models to work with 3D datasets. We also discuss the key challenges in applying diffusion models to three-dimensional (3D) vision, such as handling occlusions and varying point densities, and the computational demands of high-dimensional data. Finally, we discuss potential solutions, including improving computational efficiency, enhancing multimodal fusion, and exploring the use of large-scale pretraining for better generalization across 3D tasks. This paper serves as a foundation for future exploration and development in this rapidly evolving field.
@@ I Introduction
For our survey, we employed a comprehensive literature search strategy to ensure a thorough exploration of the field. We began by identifying key terms and phrases relevant to our topic, including “diffusion models”, “3D vision”, and related concepts such as “generative models” and “neural networks for 3D data”. We conducted our search across multiple academic databases, including IEEE Xplore  [^1] https://ieeexplore.ieee.org/Xplore/home.jsp, arXiv  [^2] https://arxiv.org/, and Google Scholar  [^3] https://scholar.google.com/, focusing on publications from the last five years to capture recent advances. Additionally, we prioritized peer-reviewed journal articles, conference papers, and preprints, ensuring the inclusion of high-quality and cutting-edge research. Using this strategy, our aim was to provide a comprehensive and up-to-date survey of diffusion models in 3D vision.
@@ IV-A Unconditional Generation
Unconditional 3D generation refers to the process of generating three-dimensional shapes or objects without being guided by any specific input or condition, such as a class label, image, or textual prompt. In this approach, the model learns to create diverse 3D structures based on patterns it has identified from training data, often starting from random noise or latent variables. This technique is used to synthesize new 3D shapes, scenes, or models in applications like 3D design, virtual environments, and gaming. The goal is to produce realistic and varied 3D outputs by capturing the underlying data distribution without external guidance.
Zhou et al.[15] introduce a new framework called Point-Voxel Diffusion, which is one of the earliest works to utilize diffusion models for generating 3D objects. This method combines denoising diffusion models with a probabilistic generative model of 3D shapes represented as point-voxel mixtures. It generates 3D shapes by performing a series of denoising steps that reverse diffuse the observed point cloud data back to Gaussian noise.
Ren et al.[16] observe that in the early stages of the diffusion process, the attention mechanism has a greater impact on the generation of the overall shape, while in the later stages, convolution operations significantly influence the local surface quality of the generated point cloud. Therefore, they propose a time-varying denoising model that combines convolution and Transformer architectures. This model rebalances global and local features by generating optimizable masks at each time step, enabling dynamic feature fusion throughout the process.
Zheng et al.[49] introduce a novel diffusion model-based pre-training method for point clouds called PointDif, aimed at enhancing the performance of point cloud models in downstream tasks such as classification, segmentation, and detection. PointDif treats the point cloud pretraining task as a conditional point-to-point generation problem and proposes a recurrent uniform sampling optimization strategy, enabling the model to recover uniformly from different noise levels.
@@
Shim et al.[13] propose a two-stage 3D shape generation framework called SDF-Diffusion. In the first stage, a low-resolution SDF of the 3D shape is generated. Then, this low-resolution SDF is used as a condition for the second stage, where a diffusion model is employed for super-resolution processing to produce a high-resolution SDF. Since this method uses SDF as the representation of the 3D shape, it allows for direct mesh reconstruction using the Marching Cubes algorithm and CNNs.
Mo et al.[17] introduce a new method called DiT-3D, which converts point clouds into voxel representations and utilizes DDPMs to perform denoising operations on the voxelized point clouds, thereby generating high-quality 3D shapes. The authors initialize DiT-3D using a pretrained DiT-2D model from ImageNet and fine-tune it to enhance performance in 3D generation tasks.
Müller et al.[50] introduce a method called DiffRF, which operates directly on the radiance field representation of 3D objects, using a 3D DDPM model to generate radiance fields directly on an explicit voxel grid. This method leverages 3D supervision and volumetric rendering to guide the generation process, enabling the unconditional synthesis of high-fidelity 3D assets.
@@
Shue et al.[51] introduce a novel 3D-aware generative model called Triplane Diffusion, designed for generating neural fields. The researchers convert 3D training scenes into a set of 2D feature planes (referred to as triplanes) and then train directly on these representations using existing 2D diffusion models to generate high-quality and diverse 3D neural fields. This approach leverages the powerful capabilities of 2D diffusion models and applies them to 3D shape generation, achieving state-of-the-art 3D generation results across multiple ShapeNet object categories.
Koo et al.[52] introduce a cascaded diffusion model called SALAD (Shape PArt-Level LAtent Diffusion Model), based on part-level implicit 3D representations, where each part is described by an independent embedding vector, which includes both external parameters (such as position and size) and internal attributes (such as geometric details). SALAD first learns diffusion in the lower-dimensional subspace (external parameters) and then performs conditional diffusion in the higher-dimensional subspace (internal attributes), enabling high-quality 3D shape generation with flexible editing capabilities.
Chou et al.[18] propose Diffusion-SDF, a method for 3D shape generation and completion. It uses SDFs as the 3D representation, where various signals (e.g., point clouds, 2D images) are parameterized by a neural network to capture their geometry. By modulating SDFs as 1D latent vectors and using these latent vectors as training data, the model learns the reverse diffusion process by predicting denoised vectors to generate samples from Gaussian noise.
Nam et al.[53] propose a method called 3D-LDM for generating neural implicit representations of 3D shapes. The authors first construct a latent space of 3D shapes based on Signed Distance Functions (SDFs) by training an auto-decoder. This auto-decoder associates each shape with a latent vector, enabling the encoding and decoding of shapes. Subsequently, 3D-LDM applies a diffusion model in the latent space of the 3D shape auto-decoder, allowing for the generation of 3D shapes with continuous surfaces.
Vahdat et al.[54] introduce a novel 3D shape generation model called LION. LION uses a Variational Autoencoder (VAE) framework that incorporates a hierarchical latent space, which includes a vector-valued latent representation for the global shape and a latent space for the point cloud structure. The core of LION involves training two levels of denoising diffusion models (DDMs) within these latent spaces to generate 3D shapes.
@@
Zhang et al.[21] introduce a new 3D shape representation method called 3DShape2VecSet, which combines concepts from Radial Basis Functions (RBF), cross-attention, and self-attention. This representation is suitable for transformer processing and can be used for neural fields and generative diffusion models. The authors also propose a two-stage model: the first stage involves an autoencoder that encodes 3D shapes into a latent space, while the second stage trains a diffusion model within the learned latent space. This approach achieves outstanding performance across various 3D shape generation and reconstruction tasks.
Lyu et al.[19] introduce a new conditional point diffusion refinement paradigm called Point Diffusion-Refinement (PDR), which is among the first to apply diffusion models to point cloud completion task. PDR first uses a DDPM to generate a coarse completed point cloud, and then applies a Refinement Network to further refine the results.
Kasten et al.[55] introduces a method called SDS-Complete, which utilizes a pre-trained text-to-image diffusion model along with the textual semantics of a given incomplete point cloud of an object to achieve a complete surface representation of the object. SDS-Complete can complete various objects at test time through optimization without the need for expensive 3D data collection, allowing the algorithm to generalize to objects that perform poorly in the training set or in real-world settings.
Nunes et al.[20] propose a LiDAR scene completion method based on DDPMs that operates directly on point cloud data rather than relying on voxel grid representations. The authors redefine the denoising scheme used in DDPMs by locally adding noise to each point instead of normalizing the input data to a noise range. This approach enables the model to learn detailed structural information of the scene effectively.
@@ IV-B Image to 3D
Image-to-3D generation focuses on converting 2D images into 3D representations. Diffusion models can help infer depth, structure, and texture from one or more 2D views, producing accurate 3D reconstructions. This task is pivotal in applications like 3D reconstruction from images or enhancing single-view 3D predictions by progressively refining the model’s understanding of spatial structure. Recent works have explored using diffusion models to tackle the challenging task of reconstructing 3D objects from single images. Diffusion models provide powerful generative priors that can help constrain this ill-posed inverse problem.
Wang et al.[56] propose a method called Score Jacobian Chaining (SJC), which applies the chain rule to multiply the gradients of the diffusion model with the Jacobian matrix of a differentiable renderer, thus aggregating 2D image gradients into 3D asset gradients. This enables the use of pretrained 2D diffusion models for 3D generation tasks without requiring additional 3D data.
Xu et al.[57] propose a CLIP-guided sampling strategy that effectively combines the prior knowledge of the diffusion model with a reference image. This method enables the reconstruction of realistic, 3D-consistent objects from a single image without requiring precise depth estimation. Additionally, they introduced scale-invariant depth supervision, reducing the need for accurate multi-view consistent depth estimation.
Szymanowicz et al.[58] propose a method called Viewset Diffusion, which uses multi-view 2D data to train a diffusion model for generating 3D objects. By integrating a bijective mapping between the viewset and the 3D model within the denoising network itself, the method ensures that the generated viewset and the 3D model are geometrically consistent.
Melas-Kyriazi et al.[59] propose RealFusion, in which they utilize the prior captured in the diffusion model to make up for the missing information. By optimizing a specific text prompt, they enabled the diffusion model to generate other possible views of a given object, reconstructing a 360° 3D model of the object from a single image. This achieved high-quality 3D reconstruction of arbitrary objects without 3D supervision.
Liu et al.[60] utilize 2D diffusion models to generate multi-view images, combined with a cost volume-based neural surface reconstruction method to reconstruct 3D meshes from multi-view predictions. This approach enables the rapid reconstruction of high-quality 3D mesh models from a single image, significantly reducing reconstruction time while more accurately capturing the details and structure of the input image.
Liu et al.[61] first fine-tunes a 2D diffusion model to predict a consistent set of images from six viewpoints, providing coherent views for 3D reconstruction. It then employs a 3D diffusion model under multi-view conditions to predict a textured mesh from the consistent multi-view images, using a coarse-to-fine approach. geBy leveraging multi-view images as supervision and applying lightweight optimization techniques, the method enhances the texture quality of the generated mesh.
Hu et al.[62] introduce EfficientDreamer, a novel method for 3D content creation that first employs a new 2D diffusion model to generate a composite image containing four orthogonal view sub-images based on a given text prompt, which provides a strong geometric structure prior and enhancing 3D consistency. Subsequently, a 3D synthesis fusion network integrates the orthogonal view diffusion model with a pretrained 2D diffusion model to address the “Janus problem” (the problem of multi-facetedness) in text-driven 3D generation, significantly improving the stability of the generation process and the quality of the 3D content.
Long et al.[63] introduce a cross-domain diffusion model for generating multi-view normal maps and corresponding color images, aimed at enhancing the geometric details and visual consistency of the generated results. It also incorporates a multi-view cross-domain attention mechanism and a geometry-aware normal fusion algorithm to facilitate information exchange between different views and modalities, thereby improving the consistency and quality of the generated outputs.
Tang et al.[64] propose a method called Make-It-3D, which leverages 2D diffusion models as 3D priors. Through a two-stage optimization pipeline, it first optimizes NeRF by combining constraints from the reference image and the diffusion prior on new views. Then, it transforms the rough model into a textured point cloud and further enhances realism using the diffusion prior, while also utilizing high-quality textures from the reference image to create high-fidelity 3D objects.
Zheng et al.[65] present a two step diffusion model-based 3D shape generation framework called LAS-Diffusion, in which the occupancy-diffusion first generates a low-resolution occupancy field to approximate the shape’s shell and then a SDF-diffusion synthesizes a high-resolution signed distance field within the occupancy voxels to extract fine geometric details. Notably, the authors employ a novel view-aware local attention mechanism that utilizes 2D image patch features to guide the learning of 3D voxel features, significantly enhancing local controllability and the model’s generalization ability.
Lin et al.[66] present Consistent123, which consists of two stages. In the first stage, it leverages 3D structural priors for extensive geometric utilization and incorporates a CLIP-based case-aware adaptive detection mechanism. The second stage introduces 2D texture priors, which gradually take a dominant guiding role to refine the details of the 3D model. Throughout the optimization process, the model emphasizes 3D structure guidance while progressively highlighting the importance of 2D priors over time to optimize texture details, achieving higher quality 3D reconstruction.
Qian et al.[67] introduce a method called Magic123 for generating high-quality 3D mesh models from a single unposed image. Magic123 employs a two-stage optimization process, first optimizing a neural radiance field (NeRF) to produce a rough geometric shape, and then using a memory-efficient differentiable mesh representation to generate a high-resolution mesh. In both stages, the generation of new views is guided by combining 2D and 3D diffusion priors, with a trade-off parameter that controls the exploration (more imaginative) and exploitation (more precise) of the generated geometry.
Tang et al.[68] distinct from previous methods reliant on NeRF and transformers, employs 3D Gaussian splatting and introduce a new 3D representation method called multi-view Gaussian features, which can be fused together for differentiable rendering. And it presents a novel framework that utilizes multi-view Gaussian features and an asymmetric U-Net backbone to quickly generate high-resolution 3D models from single-view images or text prompts.
@@
Xu et al.[69] introduce a novel feedforward 3D generation model called GRM (Gaussian Reconstruction Model) designed for efficiently reconstructing and generating 3D assets from sparse view images. At its core, GRM employs a Transformer-based network architecture that converts input pixels into 3D Gaussian values aligned with the pixels. These Gaussian values are then reprojected to form a set of densely distributed 3D Gaussian values that represent the geometry and appearance of the scene. This method not only enhances the quality and efficiency of reconstruction but also demonstrates potential in generation tasks, such as text-to-3D and image-to-3D.
Diffusion models show great promise as generative priors to address the ambiguity in reconstructing 3D objects from limited 2D supervision. Key ideas include incorporating differentiable rendering into the denoising process and using large 2D diffusion models to synthesize novel views.
Xu et al.[70] introduce InstantMesh, which aims to rapidly generate high-quality 3D mesh models from a single image while ensuring that the generated models possess good scalability for training and diverse geometric appearances. The authors combine an off-the-shelf multi-view diffusion model with a sparse-view reconstruction model based on the Large Reconstruction Model (LRM) architecture to generate consistent multi-view images, and then directly predict the 3D mesh using the LRM framework.
Liu et al.[71] propose a model called MeshDiffusion. To address the issue of irregular topology across different object categories, they utilize Deformable Tetrahedral Grids to represent 3D meshes. The model is based on Score-Based Generative Modeling, and it is trained directly on the parametric representation of 3D meshes. By training a diffusion model, MeshDiffusion generates 3D meshes effectively.

### v2

@@ Abstract
In recent years, 3D vision has become a crucial field within computer vision, powering a wide range of applications such as autonomous driving, robotics, augmented reality (AR), and medical imaging. This field relies on the accurate perception, understanding, and reconstruction of 3D scenes from 2D data sources like images and videos. Diffusion models, originally designed for 2D generative tasks, offer the potential for more flexible, probabilistic approaches that can better capture the variability and uncertainty present in real-world 3D data. However, traditional methods often struggle with efficiency and scalability. In this paper, we review the state-of-the-art approaches that leverage diffusion models for 3D visual tasks, including but not limited to 3D object generation, shape completion, point cloud reconstruction, and scene understanding. We provide an in-depth discussion of the underlying mathematical principles of diffusion models, outlining their forward and reverse processes, as well as the various architectural advancements that enable these models to work with 3D datasets. We also discuss the key challenges in applying diffusion models to 3D vision, such as handling occlusions and varying point densities, and the computational demands of high-dimensional data. Finally, we discuss potential solutions, including improving computational efficiency, enhancing multimodal fusion, and exploring the use of large-scale pretraining for better generalization across 3D tasks. This paper serves as a foundation for future exploration and development in this rapidly evolving field.
@@ I Introduction
For our survey, we employed a comprehensive literature search strategy to ensure a thorough exploration of the field. We began by identifying key terms and phrases relevant to our topic, including “diffusion models”, “3D vision”, and related concepts such as “generative models” and “neural networks for 3D data”. We conducted our search across multiple academic databases, including but not limited to IEEE Xplore  [^1] https://ieeexplore.ieee.org/Xplore/home.jsp, arXiv  [^2] https://arxiv.org/, and Google Scholar  [^3] https://scholar.google.com/, focusing on publications from the last five years to capture recent advances. Additionally, we prioritized peer-reviewed journal articles, conference papers, and preprints, ensuring the inclusion of high-quality and cutting-edge research. Using this strategy, our aim was to provide a comprehensive and up-to-date survey of diffusion models in 3D vision.
@@ II-A Introduction to Diffusion Models
In this Section, we first give an overall overview of diffusion models and their fundamental principles. Then, we introduce various core diffusion models and how they can be applied to different 3D Vision tasks.
@@ III-C3 Noise and Outliers
{forest}
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hidden-draw, rounded corners, align=left, minimum width=2em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=3.5em,font=,, where level=2text width=4.5em,font=,, where level=3text width=3.0em,font=,, [ 3D Diffusion task, ver [ Unconditional Generation, text width= 5.5em [ Explicit, text width= 1.8em [ RenderDiffusion [49], Point-Voxel Diffusion [15], TIGER [16], PointDif [50], DiT-3D [17]v PDR [19], SDS-Complete [51],
Nunes et al.[20], LION [52], 3DShape2VecSet [21] , leaf, text width=26em ] ] [ Implicit, text width= 1.8em [ SDF-Diffusion [13], Diffusion-SDF [18], DiffRF [53], Triplane Diffusion [54], 3D-LDM [55] , leaf, text width=20em ] ] ] [ Image to 3D, text width= 2.8em [ SJC [56], Neurallift-360 [57], Viewset Diffusion [58], MeshDiffusion [59], RealFusion [60], One-2-3-45 [61], EfficientDreamer [62], Wonder3d [63],
InstantMesh [64], Make-It-3D [65], LAS-Diffusion [66], Consistent123 [67], Magic123 [68], Lgm [69], GRM [70] , leaf, text width=31em ] ] [ Text to 3D, text width= 2.5em [ Point-e [71], Diffusion-SDF [72], Latent-NeRF [73], 3DFuse [74], Fantasia3D [75], STPD [76], GSGEN [77], AYG [78], GaussianDreamer [79],
ProlificDreamer [80], 4d-fy [81], Dreamtime [82], DreamFusion [83], Direct2.5 [84] , leaf, text width=30em ] ] [ Texture Generation, text width= 4.0em [ TexFusion [85], Point-UV Diffusion [86], 3DGen [87], TexOct [88], TextureDreamer [89], CRM [90], TexPainter [91], One-2-3-45++ [92], DreamMat [93],
Paint3D [94], Text2Tex [95], TEXTure [96], SceneTex [97] , leaf, text width=32em ] ] [ Human Avatar Generation, text width= 6.0em [ Avatar, text width= 1.5em [ RODIN [98], AvatarCraft [99], Chen et al.[100], DiffusionGAN3D [101], UltrAvatar [102], UV-IDM [88], , leaf, text width=22.5em ] ] [ Pose, text width= 1.5em [ PrimDiffusion [103], Chupa [104], DINAR [105], SiTH [106], ScoreHMR [107], HumanNorm [108], DreamHuman [109] , leaf, text width=25.5em ] ] [ Action, text width= 1.5em [ AnimateMe [110], InterDiff [111], PhysDiff [112] , leaf, text width=10.5em ] ] ] [ Scene Generation, text width= 4.0em [ Image-guided [ Kim et al.[113], Wynn et al.[114], LucidDreamer [115], BlockFusion [116], LiDMs [23] , leaf, text width=19em ] ] [ Text-guided [ SceneScape [117], Po et al.[118], Text2NeRF [119] , leaf, text width=11em ] ] [ Scene graph-guided [ GraphDreamer [120], DiffuScene [121], EchoScene [122] , leaf, text width=12em ] ] ] [ 3D Editing & Manipulation, text width= 6.0em [ DATID-3D [123], SKED [124], Vox-E [125], HeadSculpt [126], DreamEditor [127], SketchDream [128], GaussianEditor [129],
3D Paintbrush [12], Pandey et al.[130], APAP [131] , leaf, text width=27em ] ] [ Novel View Synthesis, text width= 4.7em [ NeRDi [132], HOLODIFFUSION [133], Tseng et al.[134], SparseFusion [135], Zero-1-to-3 [136], Chan et al.[137], Zero123++ [138], DiM [139],
NerfDiff [140], ViewDiff [141], Consistent-1-to-3 [142], SyncDreamer [143] , leaf, text width=30.5em ] ] [ Depth Estimation, text width= 4.5em [ DDP [144], DADP [145], DiffusionDepth [146], Atlantis [147], Tosi et al.[148], SEDiff [149], ECoDepth [22], DDVM [150], Marigold et al.[151] , leaf, text width=30.5em ] ] ]
Figure 2: Taxonomy of 3D Diffusion Tasks. .
@@ IV-A Unconditional Generation
Unconditional 3D generation refers to the process of generating three-dimensional shapes or objects without being guided by any specific input or condition, such as a class label, image, or textual prompt. In this approach, the model learns to create diverse 3D structures based on patterns it has identified from training data, often starting from random noise or latent variables. This technique is used to synthesize new 3D shapes, scenes, or models in applications like 3D design, virtual environments, and gaming. The goal is to produce realistic and varied 3D outputs by capturing the underlying data distribution from noises without external guidance.
Anciukevičius et al.[49] are the first to achieve 3D model generation using 2D-based DDPMs. They generate and render intermediate 3D object representations at each denoising step, introducing a strong inductive bias during the diffusion process, which can provide 3D-consistent representations with only 2D supervision. As one of the earliest works to utilize 3D diffusion models for shape generation, Zhou et al.[15] introduce a framework called Point-Voxel Diffusion, which combines denoising diffusion models with a probabilistic generative model of 3D shapes represented as point-voxel mixtures. It generates 3D shapes by performing a series of denoising steps that reverse diffuse the observed point cloud data back to Gaussian noise. To generate better 3D point cloud object, Ren et al.[16] observe that in the early diffusion stages, the attention mechanism has a greater impact on the overall shape, while in the later stages, convolution operations significantly influence the local surface quality. Therefore, they propose a time-varying denoising model that combines convolution and Transformer [46] architectures, which rebalances global and local features by generating optimizable masks at each time step, enabling dynamic feature fusion throughout the process. Zheng et al.[50] introduce a novel diffusion model-based pre-training method for point clouds called PointDif, aimed at enhancing the performance of point cloud models in downstream tasks such as classification, segmentation, and detection. PointDif treats the point cloud pretraining task as a conditional point-to-point generation problem and proposes a recurrent uniform sampling optimization strategy, enabling the model to recover uniformly from different noise levels. Different from previous point cloud-based method, Mo et al.[17] convert point clouds into voxel representations and utilizes DDPMs to perform denoising operations on the voxelized point clouds, thereby generating high-quality 3D shapes. The authors initialize their DiT-3D using a pretrained DiT-2D model [152] from ImageNet and fine-tune it to enhance performance in 3D generation tasks.
@@
In addition to directly generating 3D point clouds, some studies attempt to use DDPM to generate 3D shapes in an implicit space. Shim et al.[13] propose a two-stage 3D shape generation framework which operates on the SDFs. In the first stage, a low-resolution SDF of the 3D shape is generated. Then, this low-resolution SDF is used as a condition for the second stage, where a diffusion model is employed for super-resolution processing to produce a high-resolution SDF. Since this method uses SDF as the representation of the 3D shape, it allows for direct mesh reconstruction using the Marching Cubes algorithm and CNNs. Similarly, Chou et al.[18] propose Diffusion-SDF, which also uses SDFs as the 3D representation. By modulating SDFs as 1D latent vectors and using these latent vectors as training data, the model learns the reverse diffusion process by predicting denoised vectors to generate samples from Gaussian noise.
@@
Besides SDFs, Müller et al.[53] research DDPMs on radiance fields (RFs), which operates directly on the RFs representation of 3D objects, using a 3D DDPM model to generate radiance fields directly on an explicit voxel grid. Moreover, Shue et al.[54] introduce a novel 3D-aware generative model called Triplane Diffusion, designed for generating neural RFs. The researchers convert 3D training scenes into a set of 2D feature planes (triplanes) and then train directly on these representations using existing 2D diffusion models to generate high-quality and diverse 3D neural RFs, leveraging the powerful capabilities of 2D diffusion models.
@@
Compared to directly performing denoising on the features of 3D models, some studies [55, 52, 21] have found that first using a VAE for feature space compression, followed by applying DDPMs in the compressed space, yields better results. Nam et al.[55] propose a method called 3D-LDM for generating neural implicit representations of 3D shapes. The authors first construct a latent space of 3D shapes based on Signed Distance Functions (SDFs) by training an auto-decoder. This auto-decoder associates each shape with a latent vector, enabling the encoding and decoding of shapes. Subsequently, 3D-LDM applies a diffusion model in the latent space of the 3D shape auto-decoder, allowing for the generation of 3D shapes with continuous surfaces. Vahdat et al.[52] introduce a novel 3D shape generation model called LION. LION uses a Variational Autoencoder (VAE) framework that incorporates a hierarchical latent space, which includes a vector-valued latent representation for the global shape and a latent space for the point cloud structure. The core of LION involves training two levels of denoising diffusion models (DDMs) within these latent spaces to generate 3D shapes. Zhang et al.[21] introduce a new 3D shape representation method called 3DShape2VecSet, which combines concepts from Radial Basis Functions (RBF), cross-attention, and self-attention. This representation is suitable for transformer processing and can be used for neural fields and generative diffusion models. The authors also propose a two-stage model: the first stage involves an autoencoder that encodes 3D shapes into a latent space, while the second stage trains a diffusion model within the learned latent space. This approach achieves outstanding performance across various 3D shape generation and reconstruction tasks.
Apart from 3D generation tasks, DDPMs are also widely used for 3D completion tasks. Lyu et al.[19] introduce a new conditional point diffusion refinement paradigm called Point Diffusion-Refinement (PDR), which is among the first to apply diffusion models to point cloud completion task. PDR first uses a DDPM to generate a coarse completed point cloud, and then applies a Refinement Network to further refine the results. Kasten et al.[51] introduces a method called SDS-Complete, which utilizes a pre-trained text-to-image diffusion model along with the textual semantics of a given incomplete point cloud of an object to achieve a complete surface representation of the object. Nunes et al.[20] propose a LiDAR scene completion method based on DDPMs that operates directly on point cloud data. The authors redefine the denoising scheme used in DDPMs by locally adding noise to each point instead of normalizing the input data to a noise range, which enables the model to learn detailed structural information of the scene effectively.
@@ IV-B Image to 3D
Image-to-3D generation focuses on converting 2D images into 3D representations. Diffusion models can help infer depth, structure, and texture from one or more 2D views, producing accurate 3D reconstructions. This task is pivotal in applications like 3D reconstruction from images or enhancing single-view 3D predictions by progressively refining the model’s understanding of spatial structure.
Recent works have explored using diffusion models to tackle the challenging task of reconstructing 3D objects from single images. Diffusion models provide powerful generative priors that can help constrain this ill-posed inverse problem. Wang et al.[56] propose a method called Score Jacobian Chaining (SJC), which applies the chain rule to multiply the gradients of the diffusion model with the Jacobian matrix of a differentiable renderer, thus aggregating 2D image gradients into 3D asset gradients. This enables the use of pretrained 2D diffusion models for 3D generation tasks without requiring additional 3D data. Xu et al.[57] propose a CLIP-guided sampling strategy that effectively combines the prior knowledge of the diffusion model with a reference image. This method enables the reconstruction of realistic, 3D-consistent objects from a single image without requiring precise depth estimation. Additionally, they introduced scale-invariant depth supervision, reducing the need for accurate multi-view consistent depth estimation. Szymanowicz et al.[58] propose a method called Viewset Diffusion, which uses multi-view 2D data to train a diffusion model for generating 3D objects. By integrating a bijective mapping between the viewset and the 3D model within the denoising network itself, the method ensures that the generated viewset and the 3D model are geometrically consistent. Liu et al.[59] propose a model called MeshDiffusion. To address the issue of irregular topology across different object categories, they utilize Deformable Tetrahedral Grids to represent 3D meshes. The model is based on Score-Based Generative Modeling, and it is trained directly on the parametric representation of 3D meshes. By training a diffusion model, MeshDiffusion generates 3D meshes effectively.
When only a single-view image of an object is provided, generating a 3D model lacks information about the object’s backside. However, the generalization ability of DDPMs allows us to generate additional multi-view images to fill in this missing information, enabling more robust 3D generation. For example, Melas-Kyriazi et al.[60] propose RealFusion, in which they utilize the prior captured in the diffusion model to make up for the missing information. By optimizing a specific text prompt, they enabled the diffusion model to generate other possible views of a given object, which achieves high-quality 3D reconstruction of arbitrary objects without 3D supervision. And Liu et al.[61] utilize 2D diffusion models to generate multi-view images, combined with a cost volume-based neural surface reconstruction method to reconstruct 3D meshes from multi-view predictions. Instead of random multi-view generation, Hu et al.[62] employ a 2D diffusion model to generate a composite image containing four orthogonal view sub-images based on a given text prompt, which provides a strong geometric structure prior and enhancing 3D consistency. And for the multi-view images generated by diffusion models, Long et al.[63] introduce a multi-view cross-domain attention mechanism and a geometry-aware normal fusion algorithm to facilitate information exchange between different views and modalities, thereby improving the consistency and quality of the generated outputs. Xu et al.[64] introduce InstantMesh, which aims to rapidly generate high-quality 3D mesh models from a single image while ensuring that the generated models possess good scalability for training and diverse geometric appearances. The authors combine an off-the-shelf multi-view diffusion model with a sparse-view reconstruction model based on the Large Reconstruction Model (LRM) architecture to generate consistent multi-view images, and then directly predict the 3D mesh using the LRM framework.
Compared to end-to-end methods, two-stage approaches, although less efficient, can achieve better accuracy and are therefore widely used in image-to-3D tasks. Tang et al.[65] propose a method called Make-It-3D, which leverages 2D diffusion models as 3D priors. Through a two-stage optimization pipeline, it first optimizes NeRF by combining constraints from the reference image and the diffusion prior on new views. Then, it transforms the rough model into a textured point cloud and further enhances realism using the diffusion prior, while also utilizing high-quality textures from the reference image to create high-fidelity 3D objects. Zheng et al.[66] present a two step diffusion model-based 3D shape generation framework called LAS-Diffusion, in which the occupancy-diffusion first generates a low-resolution occupancy field to approximate the shape’s shell and then a SDF-diffusion synthesizes a high-resolution signed distance field within the occupancy voxels to extract fine geometric details. Notably, the authors employ a novel view-aware local attention mechanism that utilizes 2D image patch features to guide the learning of 3D voxel features, significantly enhancing local controllability and the model’s generalization ability. Lin et al.[67] present Consistent123, which consists of two stages. In the first stage, it leverages 3D structural priors for extensive geometric utilization and incorporates a CLIP-based case-aware adaptive detection mechanism. The second stage introduces 2D texture priors, which gradually take a dominant guiding role to refine the details of the 3D model. Throughout the optimization process, the model emphasizes 3D structure guidance while progressively highlighting the importance of 2D priors over time to optimize texture details, achieving higher quality 3D reconstruction. Qian et al.[68] introduce a method called Magic123 for generating high-quality 3D mesh models from a single unposed image. Magic123 employs a two-stage optimization process, first optimizing a neural radiance field (NeRF) to produce a rough geometric shape, and then using a memory-efficient differentiable mesh representation to generate a high-resolution mesh. In both stages, the generation of new views is guided by combining 2D and 3D diffusion priors, with a trade-off parameter that controls the exploration (more imaginative) and exploitation (more precise) of the generated geometry.
@@
Compared to NeRF, 3DGS [39] has gained increasing attention due to its superior performance in 3D representation. In 3D generation tasks, some researchers are exploring how to combine DDPMs with 3DGS to generate more exquisite 3D models. Distinct from previous methods reliant on NeRF, Tang et al.[69] employ 3D Gaussian splatting and introduce a new 3D representation method called multi-view Gaussian features, which can be fused together for differentiable rendering. They present a novel framework that utilizes multi-view Gaussian features and an asymmetric U-Net backbone to quickly generate high-resolution 3D models from single-view images. Furthermore, Xu et al.[70] introduce a novel feedforward 3D generation model called GRM (Gaussian Reconstruction Model) designed for efficiently reconstructing and generating 3D assets from sparse view images. GRM employs a Transformer-based network architecture that converts input pixels into 3D Gaussian values aligned with the pixels. These Gaussian values are then reprojected to form a set of densely distributed 3D Gaussian values that represent the geometry and appearance of the scene.